# Toward Personalized LLM-Powered Agents: Foundations, Evaluation, and Future Directions

### v1

@@ 7.3.2. User-Substitution Benchmarks
Textual and dialogue personalization benchmarks focus on evaluating how models generate content that aligns with individual user preferences, styles, and contextual knowledge. For instance, long-form generation tasks, as in LaMP (Salemi et al., 2023) and LongLaMP (Kumar et al., 2024), test whether models can adapt outputs to evolving user interests over time, while PEFT-U (Clarke et al., 2024) reformulates classification problems into annotator-specific instances to measure user-conditioned generation capabilities. Benchmarks that leverage retrieval or knowledge graphs, such as PGraphRAG (Au et al., 2025), further examine performance in sparse or cold-start scenarios. Beyond single-turn generation, multi-turn conversational benchmarks, including PersonaConvBench (Li et al., 2025e), PER-CHAT (Wu et al., 2021), LaMP-QA (Salemi and Zamani, 2025b), DPL (Qiu et al., 2025b), REGEN (Sayana et al., 2025), and KBP (Wang et al., 2023a)—assess whether models maintain coherence, adapt to a user’s conversational style, and integrate personalized knowledge consistently across dialogue. PRISM (Kirk et al., 2024) targets more complex scenarios where user values vary culturally or contextually, evaluating the reproduction of individualized subjective choices.
@@
PrefDisco (Li et al., 2025g) 10,000 General Interactive Effectiveness, Adaptivity Auto., Rule-based Discovery Accuracy, Preference Alignment, Interaction Efficiency, Correctness
@@
LongMemEval (Wu et al., ) 500 Memory Inferred Effectiveness Auto., LLM-J Accuracy, Precision, Recall, F1, Recall@K, MAP, Coherence
@@
PAL-Bench (Huang et al., 2025f) 100 Memory Inferred Effectiveness Auto., LLM-J BLEU, Win Rate
PTBench (Huang et al., 2025d) 1,083 Tool Usage Given, Inferred Effectiveness, Robustness Auto. Accuracy
@@
PSG-Agent (Wu et al., 2025b) 2,900 Multi-domain Inferred Risk Auto. Accuracy, Precision, Recall, F1-score
PENGUIN (Wu et al., 2025e) 14,000 Multi-domain Given Risk LLM-J Risk Sensitivity, Emotional Empathy, User-specific Alignment
@@
LaMP (Salemi et al., 2023) 25,095 General Inferred Effectiveness Auto. Accuracy, F1, MAE, RMSE, ROUGE
@@
PersonaConvBench (Li et al., 2025e) 111,634 Multi-domain Inferred Effectiveness Auto., LLM-J Accuracy, ROUGE, METEOR, BLEU, SBERT similarity
@@
LaMP‑QA (Salemi and Zamani, 2025b) 29,666 Q&A Given Effectiveness Auto. ROUGE‑1, ROUGE‑L, METEOR
@@ 8.1.1. Daily Dialogue Assistants
Personalized dialogue assistants primarily focus on long-term conversational continuity, where the agent must preserve user-specific context and generate preference-consistent responses across multiple interactions. These agents serve a wide range of functions, from general-purpose conversational support to open-domain recommendation and information retrieval (Zhang et al., 2025h; Huang et al., 2025a). Recent research has begun to treat long-term interaction as a benchmarked personalization problem (Pakhomov et al., 2025; Shang et al., 2025). Jones et al. (2025) further highlights how memory behaviors shape user trust and expectations, emphasizing the need for careful interface design and memory management in deployed personalized assistants. On the product side, chatbots such as ChatGPT (OpenAI, 2022), Gemini (Google, 2023), and DeepSeek (DeepSeek, 2024) have operationalized persistent personalization by allowing users to enable, inspect, and delete stored memories. In recommendation and search systems, responses are further tailored to individual needs based on ongoing user interaction (Kim et al., 2025d; Wang et al., 2024f; Huang et al., 2025e), evolving user-understanding (Wang et al., 2025b; Zhang et al., 2024b), alongside personalized decision-making models (Wang et al., 2024d; Zerhoudi and Granitzer, 2024).
@@ 8.1.2. Emotional Support Companions
These agents aim to provide sustained, empathetic assistance for users’ emotional well-being, where effectiveness depends strongly on user-contingent factors such as preferred tone, boundaries, and proactivity (Irfan et al., 2024). Personalization can be realized from both sides, where users may actively shape an agent’s persona and interaction dynamics to better match their needs (Zheng et al., 2025), while agents may adapt their behavior by explicitly modeling supportive strategies during interaction (Alotaibi and Alshahre, 2024). For example, ComPeer (Liu et al., 2024c) provides proactive peer support by detecting key moments in a dialogue and generating supportive check-ins based on conversation history and personalized strategies. MultiAgentESC (Xu et al., 2025e) enables strategy-aware support via a multi-agent pipeline that evaluates dialogue understanding, strategy selection, and response generation, providing a diverse set of personalized responses for complex emotional interactions. ARIEL (Sorino et al., 2024) takes affect-aware support further by integrating physiological signals to track emotional states and adjust responses accordingly. These research trends are reflected in companion-style products such as Replika and Inflection AI’s Pi, which focus on long-term personalization and relationship-building as core product features (Replika, ; AI, ).
@@ 8.1.3. Educational Agents
Educational agents use LLM-based systems to enhance both student and educator experiences by providing personalized guidance, adapting instruction to individual progress, and alleviating routine instructional tasks (Sharma et al., 2025; Liang and Tong, 2025). These agents can be divided into two main categories (Chu et al., 2025): (i) Personalized Pedagogical Agents, which automate pedagogical tasks for both teachers and students (Academy, 2025), and (ii) Domain-Specific Agents, which focus on specialized fields like math, music, and language learning, offering tailored support to address domain-specific challenges. Personalized pedagogical agents like EduAgent (Xu et al., 2024), LG-RAG (Cohn et al., 2025), and TeachTune (Jin et al., 2025b) adapt course materials and feedback based on learner profiles and contextual data. Domain-specific agents such as MathAgent (Yan et al., 2025b), Chat Melody (Jin et al., 2025c), Duolingo (Duolingo, 2025), and EduMAS (Li et al., 2024b) integrate domain-specific knowledge to enhance learning outcomes and reduce cognitive load, making learning more efficient and personalized.
@@ 8.2. Content Creation
Personalized content creation concerns the generation of standalone textual artifacts, where agents adapt style, structure, and framing to personalization targets rather than interactive dialogue. Typical tasks include personalized writing or summarization with preferences expressed through tone, formality, and rhetorical organization (Xu et al., 2025f; Zhang et al., 2024e; Li et al., 2025h). Personalization is evaluated at the artifact level, giving rise to two complementary paradigms: author-centric alignment with individual writing styles and audience-centric adaptation to reader knowledge and expectations (Salemi et al., 2025; Jangra et al., 2025; Novelo et al., 2025; Cai et al., 2025).
@@ 8.2.1. Author-Centric
Author-centric content creation treats personalization as a requirement of authorship rather than an optimization objective, arising in scenarios where users expect generated documents to consistently reflect their individual writing style, habitual phrasing, and subjective preferences. This need is especially prominent in real-world content creation tasks such as creative writing, professional documentation, and brand communication, where users evaluate quality at the artifact level and value stylistic continuity across independent outputs. Accordingly, personalization in this setting is defined by the ability of an agent to act as a stable authorial proxy, producing standalone texts that remain aligned with long-term user identity rather than momentary interaction context (Xu et al., 2025f; Zhang et al., 2024e; Li et al., 2025h). This application perspective is reflected in deployed systems such as Adobe Firefly, which adapts creative content generation to a user’s assets and historical editing patterns (3), Jasper Brand Voice, which operationalizes author-centric alignment for consistent brand and personal voice reproduction (68), and Notion AI, which generates documents conditioned on persistent workspace context and prior authorial conventions (146). Existing research further supports this paradigm by demonstrating that stable preference representations and document-level alignment are central to effective personalized content creation beyond conversational settings (Dey et al., 2025; Bu et al., 2025; Maram et al., 2025; Salemi et al., 2025; Song et al., 2026a; Ueda and Takayanagi, 2025).
@@ 8.2.2. Audience-Centric
Audience-centric content creation frames personalization around target readers rather than individual authors, focusing on generating textual artifacts that adapt to audience knowledge, expertise, interests, or usage contexts while preserving the intended informational content (Ning et al., 2025b; Duran and Aytekin, 2025; Zhang et al., 2024c). This paradigm is central to real-world applications such as differentiated instructional content for novice versus expert users, culturally adapted documentation, and audience-aware information presentation in web and software systems, where personalization operates at the level of audience groups rather than individual writing styles. In deployed systems, audience-centric personalization underlies large-scale marketing, communication, and outreach platforms, including social media and advertising content generation tailored to audience segments as exemplified by Predis.ai and Omneky Smart Ads (158; 149), as well as personalized sales and marketing communications targeted at specific recipient groups in systems such as SalesForge AI (176). Correspondingly, recent research emphasizes audience alignment, relevance, and comprehension as defining criteria for personalization quality in this setting, reflecting practical demands observed in audience-driven content delivery and assessment scenarios (Shi et al., 2025a; Song et al., 2026b; Tan et al., 2025b; Lyu et al., 2026; Fu et al., 2025; Salvi et al., 2025).
@@ 8.3. Delegation Assistants
Delegation assistants are personalized agents that act on behalf of users by internalizing preferences, managing information, planning actions, and supporting complex cognitive workflows. They operate over extended interaction horizons and assume delegated responsibilities aligned with user-specific goals and contexts.
@@ 8.3.1. Information Handling
Information handling concerns the long-term management of user-specific knowledge as a delegated cognitive responsibility. Delegation assistants acquire, organize, retain, and surface relevant information across tasks and over extended time horizons, reducing user cognitive load and interaction overhead. Recent work frames information access itself as a personalized service, emphasizing agentic retrieval and memory-centric designs. ARAG (Maragheh et al., 2025) models retrieval as an agent-driven process that dynamically selects and synthesizes information tailored to individual users, while SPARK (Chhetri et al., 2025) extends this paradigm to personalized search through coordinated retrieval and knowledge sharing under user context. Beyond retrieval, effective information handling requires persistent memory maintenance and selective recall. Systems such as Mr.Rec(Huang et al., 2025b) demonstrate how memory and reasoning modules jointly support personalized assistants that retain user-relevant knowledge beyond single tasks, and log-contextualized RAG (Cohn et al., 2025) shows how accumulated interaction histories enable continuity without repeated user input. These directions are reflected in deployed delegation assistants including Rewind AI (Rewind AI, Inc., 2026), Glean (Glean Technologies, Inc., 2026), You.com (You.com, Inc., 2026), Microsoft Copilot (Microsoft Corporation, 2026), and Mem (Mem, Inc., 2026), where personalization emerges through sustained ownership of the user’s informational context, enabling downstream planning and action to operate on a coherent, user-aligned knowledge substrate.
@@ 8.3.2. Task Planning
Task-planning delegation assistants aim to translate high-level user goals into executable action sequences while respecting user-specific preferences and contextual constraints. Personalization in this setting is essential for resolving trade-offs and prioritizing subtasks. Representative systems illustrate how personalization manifests in delegated planning. VAIAGE frames personalized travel planning as a collaborative multi-agent process, where specialized agents jointly coordinate subtasks under user-defined requirements (Liu et al., 2025a). Learning-based approaches further demonstrate that explicitly modeling user preferences during planning enables agents to generate plans that better align with individual goals and constraints (Xu et al., 2025b). Beyond single-task settings, collaborative language-agent frameworks have been applied to complex planning scenarios involving multiple constraints, improving flexibility and robustness in delegated workflows (Zhang et al., 2025a). Personalized task planning has also been explored across application domains, including embodied assistants that adapt planning behavior to human preferences in interactive environments (Han et al., 2025a), and social decision-making contexts such as participatory urban planning (Quan and Lee, 2025).
@@ 8.3.3. Behavior Adaptation
Behavior adaptation operates over extended time horizons, where agents align their interaction style, initiative, and decision tendencies with a user’s habitual patterns and expectations. It enables reliable delegation by internalizing stable routines while remaining responsive to evolving goals, thereby reducing user supervision and cognitive effort. Representative systems illustrate how such adaptive delegation is realized in practice. MEAgent models persistent user habits in mobile assistant settings, enabling stable delegation across daily activities (Wang et al., 2026), while memory-centric frameworks allow agents to accumulate and reuse user-specific knowledge beyond single interactions (Westhäußer et al., 2025b; Wang et al., 2025f). Test-time personalization approaches such as PersonaAgent show that adaptive behaviors can be realized dynamically by grounding responses in inferred user personas without retraining (Zhang et al., 2025c). As delegation becomes more autonomous, agents increasingly exhibit proactive behaviors, anticipating user needs rather than remaining purely reactive (Sun et al., 2025b). At the same time, empirical studies of LLM-based GUI agents caution that unconstrained behavior adaptation can lead to misalignment with human values, highlighting the need for controllable personalization in real-world delegation assistants (Gebreegziabher et al., 2026).
@@ 8.4.1. Healthcare
In healthcare, personalization is shifting from static question answering to longitudinal, stateful interaction, where effective agents must support a patient trajectory rather than a single-turn diagnosis. Frameworks such as MedChain (Liu et al., 2024a) and MedPlan (Hsu et al., 2025) organize multi-stage clinical pipelines that span history elicitation, symptom clarification, and treatment planning, emphasizing active querying to resolve uncertainty before delivering user-specific guidance. To support continuity of care, many systems maintain persistent user state, which also heightens the need for safeguards against over-accommodation and unsafe agreement (Westhäußer et al., 2025c). In response, constellation-style architectures such as Polaris and ColaCare decouple patient-facing communication from clinical reasoning, pairing an empathetic interface agent with specialist verification modules that check recommendations against guidelines (Mukherjee et al., 2024; Wang et al., 2025h; Jang et al., 2025). For data-intensive workflows, EHRAgent further reduces hallucination risk by grounding personalized EHR analysis in deterministic code execution rather than purely generative reasoning (Shi et al., 2024; Jin et al., 2024). Finally, evaluation environments such as AgentClinic (Schmidgall et al., 2024) highlight the need to stress-test personalized clinical agents with diverse simulated patient behaviors to assess robustness under realistic interaction dynamics (Lyu et al., 2025).

### v2

@@ 7.3.2. User-Substitution Benchmarks
Textual and dialogue personalization benchmarks focus on evaluating how models generate content that aligns with individual user preferences, styles, and contextual knowledge. For instance, long-form generation tasks, as in LaMP (Salemi et al., 2024) and LongLaMP (Kumar et al., 2024), test whether models can adapt outputs to evolving user interests over time, while PEFT-U (Clarke et al., 2024) reformulates classification problems into annotator-specific instances to measure user-conditioned generation capabilities. Benchmarks that leverage retrieval or knowledge graphs, such as PGraphRAG (Au et al., 2025), further examine performance in sparse or cold-start scenarios. Beyond single-turn generation, multi-turn conversational benchmarks, including PersonaConvBench (Li et al., 2025f), PER-CHAT (Wu et al., 2021), LaMP-QA (Salemi and Zamani, 2025), DPL (Qiu et al., 2025b), REGEN (Sayana et al., 2025), and KBP (Wang et al., 2023a)—assess whether models maintain coherence, adapt to a user’s conversational style, and integrate personalized knowledge consistently across dialogue. PRISM (Kirk et al., 2024) targets more complex scenarios where user values vary culturally or contextually, evaluating the reproduction of individualized subjective choices.
@@
PrefDisco (Li et al., 2025h) 10,000 General Interactive Effectiveness, Adaptivity Auto., Rule-based Discovery Accuracy, Preference Alignment, Interaction Efficiency, Correctness
@@
LongMemEval (Wu et al., 2025a) 500 Memory Inferred Effectiveness Auto., LLM-J Accuracy, Precision, Recall, F1, Recall@K, MAP, Coherence
@@
PAL-Bench (Huang et al., 2025d) 100 Memory Inferred Effectiveness Auto., LLM-J BLEU, Win Rate
PTBench (Huang et al., 2025c) 1,083 Tool Usage Given, Inferred Effectiveness, Robustness Auto. Accuracy
@@
PSG-Agent (Wu et al., 2025c) 2,900 Multi-domain Inferred Risk Auto. Accuracy, Precision, Recall, F1-score
PENGUIN (Wu et al., 2025f) 14,000 Multi-domain Given Risk LLM-J Risk Sensitivity, Emotional Empathy, User-specific Alignment
@@
LaMP (Salemi et al., 2024) 25,095 General Inferred Effectiveness Auto. Accuracy, F1, MAE, RMSE, ROUGE
@@
PersonaConvBench (Li et al., 2025f) 111,634 Multi-domain Inferred Effectiveness Auto., LLM-J Accuracy, ROUGE, METEOR, BLEU, SBERT similarity
@@
LaMP‑QA (Salemi and Zamani, 2025) 29,666 Q&A Given Effectiveness Auto. ROUGE‑1, ROUGE‑L, METEOR
@@ 8. Applications
Personalized LLM-powered agents have been applied across a wide range of settings, from conversational support and content creation to delegated assistance and expert-domain workflows. These applications differ in autonomy level, risk profile, and the primary target of personalization, but all require agents to translate user-specific information into sustained, context-sensitive behavior. We organize representative applications into four groups: conversational assistants, content creation, delegation assistants, and expert support in specific domains.
@@ 8.1. Conversational Assistant
Conversational assistants are personalized agents whose primary objective is to support users through ongoing interaction. In this setting, personalization is realized through long-term dialogue continuity, context-sensitive response generation, and adaptation to user-specific preferences, tone, and support needs. We organize representative applications in this category into daily dialogue assistants, emotional support companions, and educational agents.
@@ 8.1.1. Daily Dialogue Assistants
Personalized dialogue assistants focus on maintaining long-term conversational continuity, where agents must preserve user-specific context and produce preference-consistent responses across interactions (Zhang et al., 2025f; Huang et al., 2025a). Recent work increasingly treats long-horizon interaction itself as a benchmarked personalization problem (Pakhomov et al., 2025; Shang et al., 2025), while studies of deployed assistants highlight how memory behavior shapes user trust and expectations (Jones et al., 2025). In recommendation and search settings, personalization further depends on sustained user understanding, interaction history, and preference-aware decision making (Wang et al., 2024e; Zhang et al., 2024b; Zerhoudi and Granitzer, 2024). These trends are also reflected in deployed assistants such as ChatGPT, Gemini, and DeepSeek, which support persistent user memories and controllable personalization (OpenAI, 2022; Google, 2023; DeepSeek, 2024).
@@ 8.1.2. Emotional Support Companions
Emotional support companions provide sustained and empathetic assistance for users’ well-being, where personalization depends strongly on preferred tone, boundaries, and proactivity (Irfan et al., 2024). In this setting, personalization can arise both from user-side customization of the agent’s persona and from agent-side adaptation of supportive strategies during interaction (Zheng et al., 2025; Alotaibi and Alshahre, 2024). Representative systems include ComPeer (Liu et al., 2024c), which delivers proactive peer support based on dialogue history, MultiAgentESC (Xu et al., 2025e), which uses a strategy-aware multi-agent pipeline for emotionally complex interactions, and ARIEL (Sorino et al., 2024), which further incorporates physiological signals for affect-aware adaptation. Similar priorities are reflected in companion-style products such as Replika and Pi, where long-term personalization and relationship-building are central design features (Replika, ; AI, ).
@@ 8.1.3. Educational Agents
Educational agents personalize instruction and support by adapting guidance, feedback, and learning materials to individual users (Sharma et al., 2025; Liang and Tong, 2025). A common distinction is between personalized pedagogical agents, which support general teaching and learning workflows, and domain-specific agents, which tailor assistance to specialized subjects (Chu et al., 2025). Representative pedagogical systems such as EduAgent (Xu et al., 2024) and TeachTune (Jin et al., 2025) adapt course support based on learner profiles and contextual signals, while domain-specific agents such as MathAgent (Yan et al., 2025b) and EduMAS (Li et al., 2024a) combine personalization with specialized knowledge to improve learning effectiveness. These directions are also reflected in deployed educational systems such as Duolingo (Duolingo, 2025).
@@ 8.2. Content Creation
Personalized content creation concerns the generation of standalone textual artifacts, where agents adapt style, structure, and framing to personalization targets rather than sustaining interactive dialogue (Xu et al., 2025f). In this setting, personalization is evaluated at the artifact level, giving rise to two complementary paradigms: author-centric alignment with individual writing style and audience-centric adaptation to reader knowledge, interests, or expectations (Novelo et al., 2025).
@@ 8.2.1. Author-Centric
Author-centric content creation treats personalization as a requirement of authorship, where users expect generated documents to reflect their writing style, habitual phrasing, and long-term preferences across independent outputs (Xu et al., 2025f; Zhang et al., 2024e; Li et al., 2025i). This setting is particularly important in creative writing, professional documentation, and brand communication, where quality is judged at the artifact level rather than through dialogue. Deployed systems such as Adobe Firefly reflect this paradigm by adapting generation to persistent user assets, prior documents, or stylistic conventions (Adobe, 2023; Jasper, 2023; Notion Labs, 2023). Correspondingly, recent research highlights stable preference representations and document-level alignment as central ingredients for effective authorial personalization (Dey et al., 2025; Bu et al., 2025; Salemi et al., 2025; Ueda and Takayanagi, 2025).
@@ 8.2.2. Audience-Centric
Audience-centric content creation personalizes text for target readers rather than for the author, adapting content to audience knowledge, expertise, interests, or usage contexts while preserving the intended information (Ning et al., 2025b; Duran and Aytekin, 2025; Zhang et al., 2024c). This paradigm appears in settings such as differentiated educational materials, culturally adapted documentation, and audience-aware communication, where personalization operates at the level of reader groups or segments. In practice, it is reflected in systems such as Predis.ai and SalesForge AI, which tailor generated content to intended audience groups (Predis.ai, 2023; Salesforge, 2023). Recent research similarly emphasizes audience alignment, relevance, and comprehension as key criteria for quality in audience-conditioned generation (Shi et al., 2025a; Song et al., 2026; Tan et al., 2025b; Lyu et al., 2026; Fu et al., 2025; Salvi et al., 2025).
@@ 8.3. Delegation Assistants
Delegation assistants act on behalf of users by internalizing preferences, managing information, planning actions, and adapting behavior over extended horizons. Compared with conversational assistants, they assume a stronger degree of delegated responsibility and therefore require more persistent, user-aligned representations.
@@ 8.3.1. Information Handling
Information handling concerns the long-term acquisition, organization, retention, and retrieval of user-specific knowledge as a delegated cognitive function. In this setting, personalization lies in maintaining a coherent informational context that can support downstream planning and action without repeated user input. Representative systems include ARAG (Maragheh et al., 2025) and SPARK (Chhetri et al., 2025), which treat retrieval as a personalized, agent-driven process, as well as memory-centric assistants such as Mr.Rec (Huang et al., 2025b) and log-contextualized RAG (Cohn et al., 2025), which use persistent interaction history to support continuity across tasks. Similar ideas appear in deployed systems such as Glean and Mem (Glean Technologies, 2023; Mem, 2023).
@@ 8.3.2. Task Planning
Task-planning delegation assistants translate high-level user goals into executable action sequences while respecting user-specific preferences and contextual constraints. Personalization in this setting is crucial for resolving trade-offs and prioritizing subtasks in a user-aligned way. Representative work includes VAIAGE, which frames personalized travel planning as a collaborative multi-agent process under user requirements (Liu et al., 2025a), as well as learning-based approaches showing that explicitly modeling user preferences improves planning quality and alignment (Xu et al., 2025b). Related work also extends personalized planning to collaborative, embodied, and social decision-making settings (Zhang et al., 2025a; Han et al., 2025a; Quan and Lee, 2025).
@@ 8.3.3. Behavior Adaptation
Behavior adaptation concerns how delegation assistants align interaction style, initiative, and decision tendencies with a user’s habitual patterns over time. It supports reliable delegation by internalizing stable routines while remaining responsive to evolving goals, thereby reducing supervision and cognitive burden. MEAgent models persistent user habits in mobile assistant settings (Wang et al., 2026), while memory-centric frameworks enable agents to accumulate and reuse user-specific knowledge beyond single interactions (Westhäußer et al., 2025b; Wang et al., 2025f). Test-time personalization methods such as PersonaAgent further show that adaptive delegation can be realized dynamically through inferred user personas without retraining (Zhang et al., 2025c). At the same time, recent work highlights that more autonomous behavior adaptation also increases the need for controllable personalization and safeguards against misalignment (Sun et al., 2025b; Gebreegziabher et al., 2026).
@@ 8.4.1. Healthcare
In healthcare, personalization is shifting from single-turn question answering to longitudinal, stateful interaction, where agents must support a patient trajectory rather than only isolated diagnoses. Frameworks such as MedChain and MedPlan organize multi-stage clinical pipelines spanning history elicitation, symptom clarification, and treatment planning (Liu et al., 2024a; Hsu et al., 2025). To reduce the risks of over-accommodation and unsafe agreement, recent systems increasingly adopt modular architectures that separate patient-facing communication from clinical verification, as illustrated by Polaris and ColaCare (Mukherjee et al., 2024; Wang et al., 2025h). For data-intensive settings, EHRAgent further grounds personalized EHR analysis in deterministic code execution rather than purely generative reasoning (Shi et al., 2024; Jin et al., 2024). Evaluation environments such as AgentClinic highlight the need to stress-test personalized clinical agents under diverse simulated patient behaviors (Schmidgall et al., 2024; Lyu et al., 2025).